OpenStudy (anonymous):

What is the area of a circle with a radius of 16 cm? Use pie = 3.14 and round to the nearest tenths place if necessary. Please help, i'm not good at this =/.

OpenStudy (anonymous):

A=r^2*pi =16^2*3.14 =804

OpenStudy (anonymous):

Uhm... What? ..

OpenStudy (anonymous):

\[ A=804cm ^{2}\]

OpenStudy (anonymous):

Okay, the only option near that I got is 803.8cm2, would that be the correct answer?

OpenStudy (anonymous):

yes round to 804

OpenStudy (anonymous):

Okay, thanks, Medal for you =].

OpenStudy (anonymous):

u r we com

OpenStudy (anonymous):

the answer more exactly if you use pi instae use 3.14

OpenStudy (anonymous):

You should use pi during the calculation process. Then, u round the answer at the very end. You should get the correct answer.